what is the ratio of energy stord in the inductor to the energy stored in the capacitor when 1/3 of the total energy is stored in the capacitor of an LC circuit?
1
Expert's answer
2018-02-14T08:47:08-0500
The total energy of the LC circuit
E=E_inductor+E_capacitor
Thus
E_inductor=E-E_capacitor=E-1/3 E=2/3 E
Finally
E_inductor/E_capacitor =(2/3 E)/(1/3 E)=2/1
Answer: 2/1